# Interpret the Internal States of Recommendation Model with Sparse Autoencoders

### v2

@@ 4.5. Hyper-parameter
Amazon BPRMF BaseModel 0.0386 0.0200 0.0577 0.0248 0.2362 -0.0670 1 1 5 64
RecSAE 0.0373 0.0192 0.0563 0.0240 0.3056 0.0255 5 16 57 1009
LightGCN BaseModel 0.0426 0.0224 0.0667 0.0284 0.1519 -0.1284 0 0 2 64
RecSAE 0.0413 0.0212 0.0638 0.0268 0.4538 0.0117 8 35 100 1020
SASRec BaseModel 0.0678 0.0367 0.1046 0.0460 0.3483 0.0845 0 5 10 64
RecSAE 0.0624 0.0334 0.0962 0.0419 0.7143 0.0953 46 136 265 1004
TiMiRec BaseModel 0.0649 0.0355 0.0939 0.0427 0.3140 0.0824 0 3 10 64
RecSAE 0.0587 0.0314 0.0885 0.0389 0.6288 0.0529 22 70 187 994
MovieLens BPRMF BaseModel 0.0794 0.0382 0.1386 0.0530 0.1796 0.1075 0 0 2 64
RecSAE 0.0724 0.0355 0.1300 0.0499 0.3795 0.1088 5 18 65 929
LightGCN BaseModel 0.0771 0.0373 0.1343 0.0516 0.4324 0.0742 0 5 8 64
RecSAE 0.0746 0.0373 0.1272 0.0505 0.6225 0.1355 6 36 72 701
SASRec BaseModel 0.2008 0.1049 0.3012 0.1302 0.4445 0.0965 1 1 4 64
RecSAE 0.1868 0.0973 0.2793 0.1205 0.6980 0.1086 8 36 112 970
TiMiRec BaseModel 0.2112 0.1119 0.3105 0.1369 0.5119 0.1143 0 4 7 64
RecSAE 0.2014 0.1083 0.3007 0.1333 0.6564 0.1061 4 31 81 804
LastFM BPRMF BaseModel 0.1139 0.0702 0.1478 0.0787 0.1374 0.0244 0 1 4 64
RecSAE 0.0893 0.0544 0.1175 0.0616 0.3019 0.0162 11 38 112 1019
LightGCN BaseModel 0.1485 0.0952 0.1873 0.1049 0.1959 0.1225 0 0 3 64
RecSAE 0.1203 0.0755 0.1558 0.0844 0.4667 0.0144 7 50 108 1024
SASRec BaseModel 0.2140 0.1579 0.2446 0.1656 0.3688 0.0913 1 2 8 64
RecSAE 0.1722 0.1207 0.2043 0.1288 0.5112 0.0486 30 85 229 1024
TiMiRec BaseModel 0.2147 0.1521 0.2486 0.1606 0.2127 0.0556 0 3 11 64
RecSAE 0.1774 0.1188 0.2103 0.1271 0.5332 0.0420 35 102 222 997
5. Experiments and Results
In experiments, we aim to address the following questions: 1) What is RecSAE’s overall performance in interpreting various types of recommendation models? 2) What concepts does RecSAE interpret from different models? 3) How do RecSAE hyper-parameters influence its reconstruction and interpretation performance? 4) Are RecSAE training designs effective? 5) Can RecSAE perform targeted tuning to models?
5.1. RecSAE Overall Performance (Q1)
Table 3 presents RecSAE’s overall performance in reconstruction and interpretation across four models trained on three datasets.
The recommendation performance remains strong when the original activations are replaced with RecSAE reconstructions in the interpreted model. This confirms that the latent space effectively captures most of the information in the interpreted models at the probing positions. It is important to note that RecSAE does not affect the performances of well-trained models, as reconstructions are not intended to replace the original model activations at the probing positions during interpretation.
RecSAE identifies significantly more valid concepts than various base models. While the four recommendation models exhibit limited interpretability in their activations, RecSAE extracts latent representations that offer fine-grained interpretations with high confidence scores. Intra- and inter-concept similarity evaluations further demonstrate RecSAE’s ability to identify distinct and valid concepts. Specifically, intra-concept similarities consistently surpass those of the base models. RecSAE’s inter-concept similarity remains consistently low, indicating well-separations. While some inter-concept similarities exceed those of the original model, this may be due to base models’ limited set of concepts, with fewer than 10 confidence scores exceeding 0.8. Given the small number of concepts, a low inter-concept similarity is expected.
Table 4. Comparison with Model-level Interpretable Baseline. RecSAE on simple BPRMF yields promising performance. ** represents p¡0.05 significance between best and second best.
Dataset Model H@10 N@10 H@20 N@20
Amazon ECF_single 0.0203 0.0104 0.0341 0.0138
ECF_ensemble 0.0232 0.0115 0.0388 0.0154
DistillEmb 0.0236 0.0120 0.0386 0.0158
RecSAE 0.0373** 0.0192** 0.0563** 0.0240**
MovieLens ECF_single 0.0393 0.0183 0.0675 0.0254
ECF_ensemble 0.0393 0.0202 0.0731 0.0287
DistillEmb 0.0628 0.0285 0.1078 0.0397
RecSAE 0.0724** 0.0355** 0.1300** 0.0499**
LastFM ECF_single 0.0317 0.0277 0.0625 0.0463
ECF_ensemble 0.0741 0.0516 0.0872 0.0549
DistillEmb 0.0940 0.0577 0.1249 0.0655
RecSAE 0.0893 0.0544 0.1175 0.0616
Comparisons with baseline, ECF [^3] https://github.com/zealscott/ECF and DistillEmb [^4] https://github.com/yuan-pku/Distilling-Structured-Knowledge-into-Embeddings-for-Explainable-and-Accurate-Recommendation, are shown in Table 4. ECF’s clustering explanation module is jointly trained with recommendation models, which limits model performance, as seen in the ECF_single results. To mitigate this, ECF ensemble models with different initializations. Based on the ensemble range from the paper, we test both 3-model and 5-model ensembles. The 3-model set-up yields better performance, reported as ECF_ensemble in the table. Despite additional training costs, it still underperforms compared to RecSAE’s reconstruction with BPRMF base. DistillEmb is jointly trained with a path-based model for explanation. While it performs well on the LastFM dataset, the improvements are insignificant (p¿0.19), and RecSAE with the LightGCN base significantly outperforms it. Furthermore, RecSAE does not require this replacement during interpretation. Its probing paradigm does not interfere with well-trained models, offering both performance effectiveness and interpretability.
Table 5. Examples of Interpretable Concepts derived from SASRec model on the Amazon, MovieLens and LastFM dataset. Although some result may seem like item clustering based on titles, it is actually formed by the internal states of SASRec. This pure ID-based model, trained without semantic information, encodes human-understandable concepts within its activations.
Dataset Latent ID Item ID Item Title Activation Interpretation
Amazon 193 4582 Betty Crocker Gluten Free Brownie Mix, 16-Ounce Boxes, … 0.85 gluten-free baked goods, with a particular focus on mixes for baked goods, such as brownies, cakes, cookies, and pancake
4575 Betty Crocker Gluten Free Chocolate Chip Cookie Mix, … 0.83
5882 Bisquick Pancake and Baking Mix, Gluten-Free, … 0.82
271 7644 Happy Baby Happy Munchies Rice Cakes Apple - 1.41 oz 0.76 baby food or snacks, specifically organic and healthy options for infants and toddlers,…
5060 Happy Baby Organic Baby Food 3 Homestyle Meals,… 0.75
5057 Happy Tot Organic Baby Food, Sweet Potato, Apple, Carrot, … 0.74
Movie- Lens 29 1085 Star Trek IV: The Voyage Home (1986) (Action, Adventure, Sci-Fi) 0.57 part of a science fiction franchise, particularly the Star Trek franchise, or have a strong science fiction element
1069 Star Trek: First Contact (1996) (Action, Adventure, Sci-Fi) 0.49
1083 Star Trek: The Wrath of Khan (1982) (Action, Adventure, Sci-Fi) 0.23
617 511 Aladdin (1992) (Animation, Children’s, Comedy, Musical) 0.62 feature child-friendly , such as animation, children’s stories, or musicals, often with a strong emotional
1551 Bambi (1942) (Animation, Children’s) 0.56
1311 Anastasia (1997) (Animation, Children’s, Musical) 0.46
LastFM 260 21101 I’ll Follow The Sun - Remastered 2009, … , by The Beatles 0.65 from the Beatles, with a particular focus on songs that are considered some of their most iconic and well-known
9564 Norwegian Wood (This Bird Has Flown) …, by The Beatles 0.65
4795 Strawberry Fields Forever - Remastered 2009, …, by The Beatles 0.62
804 26791 Traust, in album Futha, by Heilung (…, dark folk, nordic folk) 0.76 strong connections to Nordic culture, particularly folk and folk-inspired music, often with a focus on traditional instruments
9656 Fehu, in album Runaljod – Yggdrasil, by Wardruna (folk, nordic folk) 0.63
9660 Solringen, in album Runaljod – Yggdrasil, by Wardruna (nordic folk) 0.60
5.2. RecSAE Interpreted Concepts (Q2)
@@ 5.2.1. Case Study
We provide examples of RecSAE concepts to show the inherent interpretability in the RecSAE latents extracted from the recommendation model. The interpretation cases for SASRec on three datasets are presented in Table 5. Notably, items that activate the same latent exhibit strong semantic similarities. In the Amazon grocery dataset, the concepts range from specific food types to particular ingredients. In MovieLens, the latents span from specific movie series to various genres. In LastFM, the interpreted concepts include different singers, albums and music types. These examples highlight the meaningful concepts captured by RecSAE within the internal states of the interpreted recommendation model.
While some results may seem to reflect item clusters based on their descriptions, these ”clusters” are actually formed by the interpreted model. These purely ID-based models encode human-understandable concepts within their activations after training exclusively on user-item interactions, without any semantic information about items or users. This underscores the remarkable ability of collaborative filtering to capture meaningful information from data. RecSAE, employing an unsupervised training approach, reveals these concepts embedded in previously opaque models.
Figure 3. Concept Survey on BPRMF (left) and SASRec (right) trained on Amazon. RecSAE interprets more diverse concepts in models with stronger recommendation performance.
@@ 5.2.2. Concept Survey
The case study above demonstrates that RecSAE successfully extracts highly interpretable latents from recommendation models, and the interpreted concepts are both human-readable and reasonable. To further examine RecSAE’s interpretability, we survey the overall interpretation situation, showing that these concepts provide valuable insights into the model’s capabilities. Figure 3 presents the word cloud for BPRMF and SASRec trained on the Amazon dataset, displaying concepts with a confidence score of 1. Notably, SASRec encodes a wide range of concepts, such as ”bean”, ”hazelnut”, ”vegan”, ”beef”, and ”microwaveable”. In this dataset, the general model BPRMF achieves a rank of 0.0200 in NDCG@10, while SASRec ranks higher at 0.0367. Meanwhile, RecSAE extracts more diverse concepts from SASRec than from BPRMF, directly revealing the differences in models’ encoding abilities.
5.2.3. Human Evaluation on Concept Relevance
To further examine RecSAE’s automatic concept construction and verification pipeline, we conduct human evaluations to assess the relevance of generated concepts with items that activate the latents. Following procedures established in previous work (Lei et al., 2023), three human experts performed four-level relevance annotations (1 irrelevant; 2 relevant; 3 contain similarities among some items; 4 contain key similarities among all items) on randomly selected 100 cases with latent confidence scores greater than 0.8. The agreement between the annotators, measured by Cohen’s kappa, is 0.3565, indicating moderate consistency. The average relevance rating is 3.8900 out of 4, reflecting a high perceived quality of the interpreted concepts. These results validate RecSAE’s automatic interpretation pipeline.
5.3. Reconstruction and Interpretation Frontier with Scale Size $s$ and Sparsity Factor $k$ (Q3)

### v3

@@ 5.2.1. Overall Performance
Rec Model Method Amazon MovieLens LastFM Yelp
Base RecSAE Base RecSAE Base RecSAE Base RecSAE
[…]
TF-IDF 1 56 28 466 0 58 14 505 0 53 27 378 9 22 28 59
BPRMF LLM 5 64 110 1009 2 64 138 1019 4 64 138 1019 5 64 63 672
Improve 4.00 0.14 2.93 1.17 - 0.10 8.86 1.02 - 0.21 4.11 1.70 -0.44 1.91 1.25 10.39
TF-IDF 0 59 49 469 7 46 42 409 0 54 67 439 9 22 83 222
LightGCN LLM 2 64 130 1020 8 64 92 701 3 64 138 1024 2 64 74 831
Improve - 0.08 1.65 1.17 0.14 0.39 1.19 0.71 - 0.19 1.06 1.33 -0.78 1.91 -0.11 2.74
TF-IDF 7 53 131 468 7 50 53 508 2 53 78 506 4 20 37 65
SASRec LLM 10 64 157 1004 4 64 111 970 8 64 111 970 1 64 173 1017
Improve 0.43 0.21 0.20 1.15 -0.43 0.28 1.09 0.91 3.00 0.21 0.42 0.92 -0.75 2.20 3.68 14.65
TF-IDF 2 55 67 482 4 45 34 485 0 56 76 506 5 52 67 324
TiMiRec LLM 10 64 193 994 7 64 81 804 11 64 227 997 3 64 141 956
Improve 4.00 0.16 1.88 1.06 0.75 0.42 1.38 0.66 - 0.14 1.99 0.97 -0.40 0.23 1.10 1.95
Table 4 presents the quantitative evaluation results of the interpreted concepts. We compare the representations from the original recommendation models with the latents extracted by RecSAE, applying two independent interpretation pipelines: a TF-IDF–based term-level statistical method and a large language model (Llama-3-8B-Instruct (AI@Meta, 2024)) capable of semantic summarizations.
Across all datasets and base recommendation models, LLMs generally yield a larger number of verified semantic concepts than the TF-IDF method. This superior performance reflects their stronger semantic understanding and ability to extract coherent concepts beyond surface-level term-frequency signals. In addition, the improvements of RecSAE over the base recommendation models are evident under both interpretation methods, demonstrating that RecSAE latents possess rich interpretability, even when evaluated using a purely term-level statistical approach.
Furthermore, sequential recommendation models (SASRec and TiMiRec) generally produce more interpretable and verifiable latents than non-sequential models (BPRMF and LightGCN). These interpretation results align with their superior recommendation accuracy performance. The interpretation results from RecSAE provide additional evidence that these models are more effective at capturing meaningful patterns from collaborative signals in purely ID-based user-item interaction logs.
Finding 2: RecSAE latents can be interpreted into verified concepts with both TF-IDF and LLM-based methods.
5.2.2. Confidence Score Distribution
The above results summarize the overall performance across four models and four datasets. To gain deeper insights into the confidence score rather than only snapshots at thresholds such as 0.8 or the entire range, we plot the distributions of latent confidence scores in Figure 2. Specifically, we visualize the distributions on the Amazon dataset across three recommendation models. The results show that the majority of latents exhibit confidence scores above 0.6. Moreover, as the base recommendation model becomes stronger, the proportion of latents with confidence scores exceeding 0.9 increases. These observations indicate that the learned latents are interpretable with substantial confidence, and that their qualities are closely aligned with the modeling capability of the underlying base recommenders.
Figure 2. Confidence Score Distribution across Interpreted Latents on BPRMF, LightGCN and SASRec on Amazon dataset.
@@ 5.2.3. Case Study
We provide examples concepts inferred by RecSAE to demonstrate both the inherent interpretability of its latents and the validity of the interpreted concepts. The full latent information and corresponding concepts of RecSAE is presented in our codebase [^3] https://github.com/Alice1998/RecSAE/tree/main/results.
@@
The interpretation cases for SASRec on four datasets are presented in Table 5. We present the latent information (latent id, corresponding activated item ids and titles) and the interpreted concepts (with TF-IDF and LLM methods for interpretation). Notably, items that are activated by the same latent exhibit strong semantic similarities. In the Amazon grocery dataset, the concepts range from specific food types to particular ingredients. In MovieLens, the latents span from specific movie series to various genres. In LastFM, the interpreted concepts include different singers, albums and music types. These cases collectively demonstrate that RecSAE effectively captures semantically meaningful concepts within the internal representations of recommendation models.
Importantly, these item groups that activated the same latent emerge solely from the internal representations learned by the recommendation models and are extracted by RecSAE through unsupervised training. The above process does not incorporate any semantic information. It is particularly noteworthy that such interpretable concepts arise within ID-based recommendation models trained purely on collaborative user–item interaction data. This highlights the capacity of collaborative filtering methods to uncover structured semantic patterns from interaction signals alone. RecSAE, through its unsupervised framework, sheds light on these latent concepts, revealing the rich, human-understandable structures embedded within otherwise opaque recommendation models.
Table 5. Examples of RecSAE Latent information and Corresponding Concepts derived from SASRec model on four datasets. The pure ID-based model, trained without semantic information, encodes human-understandable concepts within its representations. The TF-IDF concepts are the Top-3 terms. The full logs of all latents are presented in our codebase.
Dataset -LID IID Item Title TF-IDF Concept LLM Concept
A-193 4582 Betty Crocker Gluten Free Brownie Mix,… Gluten, Free, Mix gluten-free baked goods, with a particular focus on mixes for baked goods
4575 Betty Crocker Gluten Free Chocolate Chip Cookie Mix, …
5882 Bisquick Pancake and Baking Mix, Gluten-Free, …
A-271 7644 Happy Baby Happy Munchies Rice Cakes Apple … Baby, Happy, Cake baby food or snacks, specifically organic and health options for infants
5060 Happy Baby Organic Baby Food 3 Homestyle Meals,…
5057 Happy Tot Organic Baby Food, Sweet Potato,…
A-275 7890 Decaffeinated Coffee … Keurig K-Cup Brewers Keurig, Coffee, Brewers coffee-based, particularly single-serve cups with a focus on decaf
7968 The Organic Coffee Co. Decaf Gorilla Coffee …
M-204 2496 Blast from the Past (1999) (Comedy, Romance) Romance, Seattle, Sleepless romantic comedies, featuring light-hearted, humorous storylines
2671 Notting Hill (1999) (Comedy, Romance)
539 Sleepless in Seattle (1993) (Comedy, Romance)
M-612 1376 Star Trek IV: The Voyage Home (Action, Adventure, Sci-Fi) Trek, Sci-Fi, Predator the Star Trek franchise, or have a strong science fiction element
1356 Star Trek: First Contact (Action, Adventure, Sci-Fi)
1374 Star Trek: The Wrath of Khan (Action, Adventure, Sci-Fi)
M-684 588 Aladdin (1992) (Animation, Children’s, Comedy, Musical) Animation, Aladdin, Children feature child-friendly, such as animation, children stories, or musicals
2294 Antz (1998) (Animation, Children’s)
3034 Robin Hood (1973) (Animation, Children’s)
L-254 459 Adios, in album HUSH, by EVERGLOW (pop, k-pop …) k-pop, merry, room high-energy, upbeat K-pop music with strong electronic and danceable elements …
26336 Snapping, in album Flourishing, by CHUNG HA (… k-pop)
25774 FANCY, in album FANCY YOU, by TWICE (k-pop)
L-260 21101 I’ll Follow The Sun …, by The Beatles Beatles, Remastered, Rock from the Beatles, with songs considered their most iconic and well-known
9564 Norwegian Wood (This Bird Has Flown) …, by The Beatles
4795 Strawberry Fields Forever …, …, by The Beatles
L-804 26791 Traust,… Futha, by Heilung (… dark folk, nordic folk) Yggdrasil, Runaljod, Nordic connect to Nordic culture, particularly folk music with traditional instruments
9656 Fehu,… Runaljod–Yggdrasil, by Wardruna (folk, nordic folk)
9660 Solringen,… Runaljod–Yggdrasil, by Wardruna (nordic folk)
Y-84 1372 Bartaco, in Nashville, TN, (Mexican, Nightlife, Bars) Chicken, Nashville, TN Southern-style restaurants and bars that serve traditional American comfort food
2674 Hattie B’s Hot Chicken, in Nashville, TN (American, Southern)
13966 The Stillery, in Nashville, TN (American)
Y-352 3778 Café Du Monde, in New Orleans, LA (Cafes, Restaurants) Shops, Orleans, LA Coffee shops that serve traditional New Orleans-style coffee…
4391 Dong Phuong, in New Orleans, LA (Breakfast, Brunch)
11395 Bearcat Cafe, in New Orleans, LA (Coffee, Breakfast, Brunch)
Y-846 10553 Talula’s Garden, in Philadelphia, PA (Restaurants) Bars, Parc, Philadelphia Fine dining restaurants that serve French-inspired cuisine with a focus on wine
10641 Parc, in Philadelphia, PA (French, Wine Bars, Nightlife)
13447 Suraya, in Philadelphia, PA (Beer Gardens, Nightlife)
Finding 3: RecSAE interprets human-readable concepts from ID-based recommendation models trained without any semantic information.
Figure 3. RecSAE Latent Activation Distribution of SASRec. Cases are categorized as negative (blue) or positive (orange) based on the presence of the corresponding concepts. Results show that positive cases exhibit higher activations in the corresponding latents.
We further present the randomly selected cases on latent activation distributions across datasets, as shown in Figure 3. The distributions are grouped based on whether the corresponding interpreted concept is present in the activated items. In positive cases, where the concept is present, the associated latent exhibits significantly higher activation values compared to negative cases, where the concept is absent. This clear separation indicates that RecSAE successfully aligns latent activations with the presence of corresponding concepts.
These results provide empirical evidence of both the necessity and sufficiency of latent activations for concept presence. Latents are strongly activated when their associated concepts appear (sufficiency), and are minimally activated otherwise (necessity). This further validates the interpretability and semantic consistency of RecSAE interpretations on recommendation models.
Finding 4: The presence of concepts in recommendation model outputs corresponds to distinct patterns of associated latent activations.
@@ 5.2.4. Concept Survey
The preceding case study demonstrates that RecSAE is capable of extracting highly interpretable latent concepts from recommendation models. These concepts are human-readable, semantically coherent and easy to verify. To further assess the interpretability provided by RecSAE, we conduct a comprehensive survey of the interpreted concepts, illustrating their potential to reveal meaningful insights into the model’s capabilities.
Figure 4. Concept Survey on BPRMF (left) and SASRec (right) trained on the four datasets. RecSAE with the LLM interpretation gains more diverse concepts in models with stronger recommendation performance.
We perform both qualitative and quantitative analyses of the interpreted concepts. Figure LABEL:fig:word_cloud presents word clouds generated from the latent concepts of BPRMF and SASRec across four datasets. Only concepts with a confidence score of 1 are visualized to ensure high reliability. Notably, SASRec captures a broader range of meaningful concepts across all datasets. In the Amazon (Grocery and Gourmet Food) dataset, RecSAE extracts food-related terms such as dietary restrictions (”gluten”, ”glutino”), types (”cappuccino”, ”tea”, ”pasta”, ”cheeseburger”), packaging/usage features (”microwavable”), and brand names (”Crunchmaster”). In MovieLens, genre-related tags such as ”horror”, ”musical”, ”children”, ”comedy”, ”western”, ”adventure”, as well as specific movie references like ”predator”, are identified. In LastFM, musical genres (”metal”, ”rock”, ”rap”, ”pop”), artist names (”Taylor”, ”Eilish”, ”Beatles”), album titles (”Jackboys”), and descriptive terms from song titles (e.g., ”Remastered”) are effectively captured by RecSAE. In Yelp, business types (”bar”, ”restaurant”), food types (”seafood”, ”wine”, ”brunch”, ”desserts”, ”barbecue”) and style (”American”, ”French”, ”Mexican”) are effectively captured by RecSAE.
Although BPRMF also reveals some interpretable concepts, the breadth and diversity are noticeably more limited compared to the sequential model SASRec. In the Amazon dataset, BPRMF achieves an NDCG@10 of 0.0200, while SASRec attains a higher score of 0.0367. This performance discrepancy is consistent with the richness and coverage of RecSAE’s interpreted concepts, highlighting the differences in representational capacity between the two models.
Table 6. We quantify the presence of various features (Genre in MovieLens; Music Tag, Astist, Album in LastFM) across interpreted concepts by TF-IDF method with verification score $\geq$ 0.8. Specifically, we report the number of features that appear in the concepts and the coverage of features.
Model MovieLens LastFM
Genre Single-term Tag All Tag Artist Album
BPRMF 17 (94%) 31 (15%) 61 (11%) 35 (6‰) 125 (10‰)
SASRec 16 (89%) 55 (27%) 109 (20%) 60 (10‰) 182 (14‰)
Beyond the qualitative analysis, we further quantify the coverage of RecSAE’s concepts over the item features provided in the datasets, as reported in Table 6. Since the Amazon Grocery and Gourmet Food dataset lacks distinct category labels (items are all labeled as ”Grocery and Gourmet Food”), we focus the coverage analysis on the MovieLens (genre) and LastFM (tag, artist, album) datasets. Feature coverage is measured by the ratio of the exact string match between item features and the extracted concept terms to the total number of item features. In MovieLens, all genre features are successfully identified in SASRec’s concepts. In LastFM, which contains a wider variety of feature types, SASRec’s concepts cover over 20% of item tags, in contrast to BPRMF, which covers no more than 12%. This analysis also demonstrates that RecSAE’s interpretation results correlate with the model’s encoding abilities: higher model accuracy is associated with a greater diversity of interpreted concepts.
Finding 5: RecSAE can serve as an effective tool for understanding and comparing model encoding behaviors.
5.3. Concept Quality across Activation Level
In this section, we further analyze concept quality across different activation levels. We evaluate the concept hit ratio within recommended items at varying latent activation levels, where the hit ratio measures the presence of the corresponding concept in the top-1 recommended item. The results are presented in Figure 5.
The results demonstrate a clear trend: higher activation levels correspond to a stronger presence of the associated concepts in the recommended items (blue line). Apart from the changes from level 9 to 10 on BPRMF and from level 8 to 9 on SASRec, the overall progression remains smooth, with no distinct cutoff where the concepts become indistinct as the latents lose discriminative power. Moreover, the average global hit ratios across all features are 0.016 for BPRMF and 0.043 for SASRec (red lines), which are substantially lower than the hit ratios observed when the corresponding latents are activated (blue lines). This pattern further confirms the strong associations between the latents and their respective concepts.
Figure 5. Concept Quality across Latent Activation Levels on the MovieLens Dataset, BPRMF (left) and SASRec (right). When the latent is activated, the corresponding concept hit ratio in the top-1 recommended item (blue line) is substantially higher than the global average (red line).
5.4. Human Evaluation of Interpreted Concepts
To Further assess the quality of RecSAE interpretation pipeline, we conduct a human evaluation to determine the relevance of the generated concepts to the items activating the corresponding latents. Following the annotation protocol established in RecExplainer (Lei et al., 2023), three recommendation experts independently annotated 100 randomly selected cases with a latent confidence score above 0.8. A four-level relevance scale was used: 1 - irrelevant, 2 - weakly relevant, 3 - partially similar (some items share similarities), and 4 - highly similar (all items share key similarities). Specifically, we give annotators the items activating the latent and the corresponding concepts of the latents.
The annotation results show high relevance between latents and corresponding interpreted concepts. We first assess the inter-annotator agreement to evaluate annotation quality. Following prior work (Lei et al., 2023), we compute the pairwise Cohen’s kappa between every two annotators and report the averaged value, as well as the overall Fleiss’ kappa for multi-annotator agreement. The agreement scores are 0.3565 for Cohen’s kappa and 0.3407 for Fleiss’ kappa on the 4-level annotations. These results demonstrate a fair degree of consistency among annotators, comparable to that reported in previous studies on recommendation explanation (e.g., Cohen’s kappa of 0.36 in RecExplainer (Lei et al., 2023)). The average relevance rating across all annotated cases is 3.89 out of 4, indicating that the majority of the interpreted concepts are perceived as highly coherent and semantically meaningful. These findings support the validity of RecSAE’s automated interpretation framework.
Finding 6: RecSAE interpreted concepts are perceived as relevant (3.89/4) to the items activating the corresponding latents by human experts.
5.5. Comparison with Interpretation Baseline
Figure 6. The visualization of SASRec’s attention weights on MovieLens (Kang and McAuley, 2018). It demonstrates that the attention mechanism can capture items sharing common features. However, it is inherently constrained by the feature set present in the current input sequence, making it difficult to tune the model toward features absent from the current context.
As few methods exist for interpreting different types of recommendation models, we compare the proposed RecSAE framework with baselines specific to different model families. In particular, we evaluate the attention-based explanation of SASRec. Following the original SASRec paper, which uses the same model and dataset, we refer to its attention visualization for comparison. Two disjoint sets of 200 movies are randomly selected from four categories—Science Fiction, Romance, Animation, and Horror—with the first set used as queries and the second as keys. Figure 6 illustrates the average attention weights across items, reflecting the model’s ability to capture feature-level similarities. However, a fundamental limitation of attention-based interpretation and tuning remains: while attention can distinguish feature-level similarities within the observed input, it lacks the capacity to explicitly tune the model’s preference toward unobserved target features. This motivates the need to explicitly model neurons or embeddings specific to each feature beyond attention.
6. Reconstruction Performance
In the previous section, we demonstrated the strong interpretability achieved by the RecSAE framework. In this section, we further assess its reconstruction performance to verify that the learned latent space preserves the information contained in the original model’s internal representations.
Table 7 presents the reconstruction performance. We assess recommendation accuracy using Hit Rate and NDCG by replacing the original model’s internal representations with RecSAE’s reconstructed activations at the designated probing positions. The results indicate that recommendation performance remains largely stable across all four recommendation models and four datasets, confirming that RecSAE’s latent space effectively preserves the essential information from the original representations. In comparison to a standard autoencoder (AE) with the same number of latent activations (8), RecSAE’s sparse autoencoder demonstrates significantly better reconstruction performance across all datasets. This highlights the advantage of the sparse autoencoder architecture in retaining critical information from the original model while also encouraging the formation of disentangled, mono-semantic latents that support effective interpretation.
It is worth emphasizing that RecSAE does not interfere with the performance of the base recommendation models, as its reconstructions are not used to replace the original representations during the interpretation process. The RecSAE latents are solely employed for interpretation, and the above reconstruction results show that these latents effectively preserve the information encoded in the original model representations.
Table 7. RecSAE Reconstruction Performance. We evaluate the training with downstream performance when replacing the model’s original representations with reconstructions. ”N” is NDCG. RecSAE reconstruction performances remain stable, confirming RecSAE latents effectively preserve the essential information from the original model representations.
Rec Model Method Amazon MovieLens LastFM Yelp
N@10 N@20 N@10 N@20 N@10 N@20 N@10 N@20
Base 0.0200 0.0248 0.0382 0.0530 0.0702 0.0787 0.0036 0.0063
BPRMF AE 0.0157 0.0207 0.0313 0.0429 0.0129 0.0166 0.0032 0.0061
RecSAE 0.0192 0.0240 0.0355 0.0499 0.0544 0.0616 0.0034 0.0063
Base 0.0224 0.0284 0.0373 0.0516 0.0952 0.1049 0.0069 0.0106
LightGCN AE 0.0157 0.0204 0.0320 0.0444 0.0127 0.0159 0.0044 0.0066
RecSAE 0.0212 0.0268 0.0373 0.0505 0.0755 0.0844 0.0065 0.0096
Base 0.0367 0.0460 0.1049 0.1302 0.1579 0.1656 0.0043 0.0065
SASRec AE 0.0312 0.0391 0.0441 0.0584 0.0111 0.0146 0.0040 0.0062
RecSAE 0.0334 0.0419 0.0973 0.1205 0.1207 0.1288 0.0040 0.0060
Base 0.0355 0.0427 0.1119 0.1369 0.1521 0.1606 0.0149 0.0189
TiMiRec AE 0.0275 0.0342 0.0490 0.0649 0.0150 0.0195 0.0054 0.0076
RecSAE 0.0314 0.0389 0.1083 0.1333 0.1188 0.1271 0.0119 0.0160
7. Parameter Sensitivity on Reconstruction and Interpretation
We conduct sensitivity experiments to examine the influence of hyper-parameters in the RecSAE framework. Specifically, we evaluate how the scale size $s$, which determines the dimensionality of the latent space, and the sparsity factor $k$, which controls the number of active latents per instance, affect both the reconstruction quality and the interpretability of the learned representations.
7.1. Scale Size $s$